2x-3y+z=7
x+2y-3z=7
-x+4y+3z=5

solve in three variables using cramer's rule
thank you

Answer by checkley71(8403)

2 -3 1 |
1 2 -3 | =2(6+12)+3(3-3)+(4+2)=36+0+6=42
-1 4 3 |
7 -3 1 |
7 2 -3 |=7(6+12)+3(21+15)+28-10)=126+108+18=252
5 4 3 |
2 7 1 |
1 7 -3 |=2(21+15)-7(3-3)+(5=7)=72+0+12=84
-1 5 3 |
2 -3 7 |
1 2 7 |=2(10-28)+3(5+7)+7(4+2)=-36+36+42=42
-1 4 5 |
THUS THE FOLLOWING SOLUTIONS:
X=252/42 OR X=6
Y=84/42 OR Y=2
Z=42/42 OR Z=1